# The Hilbert-space structure of free fermions in disguise

### v2

@@ 4.6 $M=6k+5$
4.7 Factoring out the symmetry: mapping to another FFD chain
@@
where on each macro-site $X,Z$, and $X^{\prime},Z^{\prime}$ refer respectively to the action on the first or second spin. Depending on the value of $M({\rm mod}~3)$, the mapping (39c) may need to be adapted for the rightmost operators $h_{m}$, but we will not discuss this in detail here. In summary, the projection has allowed us to map the model onto a chain of $\sim 2M/3$ spins, with the same energy spectrum as the original model but less degeneracies.
Going further, we could “factor out” the rest of the symmetry algebra by fixing the eigenvalues of the remaining generators $\widetilde{h}_{6l+2}\widetilde{h}_{6l+3}$. The resulting model has a Hilbert space of dimension $\sim 2^{M/2}$ and we expect that it coincides with the two-chain model presented in Ref [19] and worked out in more detail in Ref.  [27] (in that reference, “minimal representations" of the FFD algebra were presented for any value of $M$). Since this direction is beyond the scope of this work, we will not pursue it further, but it remains an interesting question for future work.
5 Ancillary free fermions in disguise
@@
5.1 The (super)symmetry algebra
The construction of the algebra $\mathcal{A}_{F^{\prime}}$ follows closely that of the extended SUSY algebra described in Ref. [26], up to subtle differences which we will highlight below.
@@ 5.1 The (super)symmetry algebra
$\chi_{m}=a_{m}~\chi^{\prime}\prod_{i=1}^{m-1}h_{i}\,,\qquad m=1,\ldots M+1\,$, (40)
@@
The same relations appeared in Ref. [19], albeit for different operators, as we now discuss. By expanding Eq. (40) in terms of Pauli matrices, we see that the operators $\chi_{m}$ defined here contain Pauli strings of the form $XYXX\ldots$. This is different from the definition of the operators denoted by $\chi_{m}$ in Ref. [19], which contain Pauli strings of the form $XXX\ldots$. This difference has crucial implications when studying the commutation relations with the physical fermions, as we will see shortly.
Following Ref. [19], it is not hard to check that the linear combinations
@@
Next, following once again Ref. [19], we can define a whole hierarchy of operators $O^{(n)}$, $E^{(n)}$ (with $O^{(1)}=O$, $E^{(1)}=E$), all commuting with $H$ and the higher conserved charges. Using the above notation,
@@
We have therefore reached our first step in identifying the symmetry algebra $\mathcal{A}_{F^{\prime}}$: it contains the operators $O^{(n)}$ for $M$ odd, and the operators $E^{(n)}$ for $M$ even. More precisely, as it will be clear from our subsequent analysis, the symmetry algebra $\mathcal{A}_{F^{\prime}}$ is generated by the $O^{(2n)}$ (resp. $E^{(2n)}$) together with the bilinears in the $O^{(2n+1)}$ (resp. $E^{(2n+1)}$), such that the extra boundary operator $\chi^{\prime}$ drops out in such operators (note that, since $\chi^{\prime}$ does not commute with all $\widetilde{h}_{m}$ operators, it acts non-trivially in $\mathcal{H}_{\widetilde{D}}$). This picture is different from the situation in Ref. [19], where the operators $E^{(n)}$ and $O^{(n)}$ all have non-trivial commutation relations with the fermions.
5.2 Commuting transfer matrices
@@
$\mathcal{E}_{2m}(u)$ $=\mathcal{E}_{2m-2}(u)+u~\mathcal{O}_{2m-2}(u)\chi_{2m}\,$, (46a)
$\mathcal{O}_{2m}(u)$ $=\mathcal{O}_{2m-2}(u)+u~\mathcal{E}_{2m-4}(u)\chi_{2m-1}\,$, (46b)
$\mathcal{O}_{2m+1}(u)$ $=\mathcal{O}_{2m-1}(u)+u~\mathcal{E}_{2m-1}(u)\chi_{2m+1}\,$, (46c)
$\mathcal{E}_{2m+1}(u)$ $=\mathcal{E}_{2m-1}(u)+u~\mathcal{O}_{2m-3}(u)\chi_{2m}\,.$ (46d)
@@
5.3 The ancillary fermions
We now have all the ingredients for diagonalizing the operators $O^{(n)}$ and $E^{(n)}$. As we will see, the diagonalization follows very closely Fendley’s original solution of the FFD Hamiltonian [19], and involves another set of ancillary free fermions commuting with the physical ones.
@@ 5.3 The ancillary fermions
$\mathbb{E}_{2m}^{\pm}(u)$ $=\mathbb{E}_{2m-2}^{\pm}(u)+u~\mathbb{O}_{2m-2}^{\mp}(u)a_{2m}^{2}$ (50a)
$\mathbb{O}_{2m}^{\mp}(u)$ $=\mathbb{O}_{2m-2}^{\mp}(u)+u~\mathbb{E}_{2m-4}^{\pm}(u)a_{2m-1}^{2}$ (50b)
$\mathbb{O}_{2m+1}^{\pm}(u)$ $=\mathbb{O}_{2m-1}^{\pm}(u)+u~\mathbb{E}_{2m-1}^{\mp}(u)a_{2m+1}^{2}$ (50c)
$\mathbb{E}_{2m+1}^{\mp}(u)$ $=\mathbb{E}_{2m-1}^{\mp}(u)+u~\mathbb{O}_{2m-3}^{\pm}(u)a_{2m}^{2}\,.$ (50d)
@@
[…]
@@
Going further, we can construct the corresponding fermionic creation and annihilation operators. Since the derivation is similar to those of Ref. [19], we omit it and only provide the final result: the ancillary creation and annihilation operators read
$\Psi^{\prime}_{\pm k}\propto\begin{cases}\mathcal{E}_{M}(\sqrt{\pm u_{k}})~E~\mathcal{E}_{M}(\sqrt{\mp u_{k}})\qquad\mbox{for $M$ even}\\ \mathcal{O}_{M}(\sqrt{\pm u_{k}})~O~\mathcal{O}_{M}(\sqrt{\mp u_{k}})\qquad\mbox{for $M$ odd}\end{cases}\,$, (52)
@@
5.4 The symmetry algebra $\mathcal{A}_{F^{\prime}}$ in terms of the ancillary fermions
@@
We are therefore ready to state our final proposition, namely that the algebra $\mathcal{A}_{F^{\prime}}$ is generated by all bilinears in the ancillary fermions $\Psi^{\prime}_{k}$. This is confirmed by a counting of dimensions presented in Table 1, cf. Sec. 3, where it is checked that the ancillary fermionic modes, together with the algebra $\mathcal{A}_{\widetilde{D}}$ of Sec. 4, indeed accounts for all the degeneracies observed in the spectrum.
5.5 The remaining zero mode for $M=6k+3$
As mentioned above, for the special case $M=6k+3$, the algebras $\mathcal{A}_{\widetilde{D}}$ and $\mathcal{A}_{F^{\prime}}$ are enough to lift all degeneracy factors of the spectrum, but one. This can be explained by the existence of an additional zero mode, which commutes both with the physical fermions and with $\mathcal{A}_{\widetilde{D}}$ and $\mathcal{A}_{F^{\prime}}$. For $M=3$, such an operator can easily be found as $(h_{1}+h_{2})\chi_{M}$. For $M=9$, it can be found as a bilinear in the $\{h_{i}\}_{i\leq M-1}$ multiplied by $\chi_{M}$. For general $M=6k+3$ we conjecture that an operator of such a form should exist, namely some multilinear combination of $\{h_{i}\}_{i\leq M-1}$ multiplied by $\chi_{M}$. Furthermore, it should be possible to relate the existence of this operator to the zero modes evidenced in [29] more generally for all $M\notin 3\mathbb{N}+1$. However, a difference between our approach and the zero modes of [29] is that the latter do not necessarily commute with the rest of the symmetry algebra, that is, with $\mathcal{A}_{\widetilde{D}}$ and $\mathcal{A}_{F^{\prime}}$.
6 Outlook
@@
Our work opens several natural directions. Most naturally, our results provide new tools for the computation of correlations functions, both in an out of equilibrium. Previously, this problem was tackled by seeking a solution to the so-called “inverse problem”, consisting in expressing local spin operators in terms of fermionic ones. In particular Ref. [29] found a few special local spin operators that admit a simple representation in terms of the free fermionic operators, making it possible to study their dynamics explicitly. In our work, we have completely characterized the set of operators generating the symmetry algebra, which is supported on the degenerate subspace and, by definition, not evolving under the FFD Hamiltonian dynamics. This allows us go further, and consider spin operators admitting a simple representation not only in terms of the fermionic ones, but also of the elements of the symmetry algebra. Finally, another natural direction pertains to extending our constructions to different FFD models, including the one presented in Ref. [26], the parafermionic generalizations developed in Refs. [24, 25], and the FFD quantum circuits studied in Ref. [27]. These directions will be explored in future research [60].

### v4

@@ IV.6 $M=6k+5$
IV.7 Factoring out the symmetry: mapping to another FFD chain
@@
where on each macro-site $X,Z$, and $X^{\prime},Z^{\prime}$ refer respectively to the action on the first or second spin. Depending on the value of $M({\rm mod}\penalty 10000\ 3)$, the mapping (39c) may need to be adapted for the rightmost operators $h_{m}$, but we will not discuss this in detail here. In summary, the projection has allowed us to map the model onto a chain of $\sim 2M/3$ spins, with the same energy spectrum as the original model but less degeneracies.
Going further, we could “factor out” the rest of the symmetry algebra by fixing the eigenvalues of the remaining generators $\widetilde{h}_{6l+2}\widetilde{h}_{6l+3}$. The resulting model has a Hilbert space of dimension $\sim 2^{M/2}$ and we expect that it coincides with the two-chain model presented in Ref [fendley2019free] and worked out in more detail in Ref. [fukai2025quantum] (in that reference, “minimal representations” of the FFD algebra were presented for any value of $M$). Since this direction is beyond the scope of this work, we will not pursue it further, but it remains an interesting question for future work.
V Ancillary free fermions in disguise
@@
V.1 The (super)symmetry algebra
The construction of the algebra $\mathcal{A}_{F^{\prime}}$ follows closely that of the extended SUSY algebra described in Ref. [fendley2024free], up to subtle differences which we will highlight below.
@@ V.1 The (super)symmetry algebra
$\chi_{m}=a_{m}\penalty 10000\ \chi^{\prime}\prod_{i=1}^{m-1}h_{i}\,,\qquad m=1,\ldots M+1\,$, (40)
@@
The same relations appeared in Ref. [fendley2019free], albeit for different operators, as we now discuss. By expanding Eq. (40) in terms of Pauli matrices, we see that the operators $\chi_{m}$ defined here contain Pauli strings of the form $XYXX\ldots$. This is different from the definition of the operators denoted by $\chi_{m}$ in Ref. [fendley2019free], which contain Pauli strings of the form $XXX\ldots$. This difference has crucial implications when studying the commutation relations with the physical fermions, as we will see shortly.
Following Ref. [fendley2019free], it is not hard to check that the linear combinations
@@
Next, following once again Ref. [fendley2019free], we can define a whole hierarchy of operators $O^{(n)}$, $E^{(n)}$ (with $O^{(1)}=O$, $E^{(1)}=E$), all commuting with $H$ and the higher conserved charges. Using the above notation,
@@
We have therefore reached our first step in identifying the symmetry algebra $\mathcal{A}_{F^{\prime}}$: it contains the operators $O^{(n)}$ for $M$ odd, and the operators $E^{(n)}$ for $M$ even. More precisely, as it will be clear from our subsequent analysis, the symmetry algebra $\mathcal{A}_{F^{\prime}}$ is generated by the $O^{(2n)}$ (resp. $E^{(2n)}$) together with the bilinears in the $O^{(2n+1)}$ (resp. $E^{(2n+1)}$), such that the extra boundary operator $\chi^{\prime}$ drops out in such operators (note that, since $\chi^{\prime}$ does not commute with all $\widetilde{h}_{m}$ operators, it acts non-trivially in $\mathcal{H}_{\widetilde{D}}$). This picture is different from the situation in Ref. [fendley2019free], where the operators $E^{(n)}$ and $O^{(n)}$ all have non-trivial commutation relations with the fermions.
V.2 Commuting transfer matrices
@@
$\mathcal{E}_{2m}(u)$ $=\mathcal{E}_{2m-2}(u)+u\penalty 10000\ \mathcal{O}_{2m-2}(u)\chi_{2m}\,$, (46a)
$\mathcal{O}_{2m}(u)$ $=\mathcal{O}_{2m-2}(u)+u\penalty 10000\ \mathcal{E}_{2m-4}(u)\chi_{2m-1}\,$, (46b)
$\mathcal{O}_{2m+1}(u)$ $=\mathcal{O}_{2m-1}(u)+u\penalty 10000\ \mathcal{E}_{2m-1}(u)\chi_{2m+1}\,$, (46c)
$\mathcal{E}_{2m+1}(u)$ $=\mathcal{E}_{2m-1}(u)+u\penalty 10000\ \mathcal{O}_{2m-3}(u)\chi_{2m}\,.$ (46d)
@@
V.3 The ancillary fermions
We now have all the ingredients for diagonalizing the operators $O^{(n)}$ and $E^{(n)}$. As we will see, the diagonalization follows very closely Fendley’s original solution of the FFD Hamiltonian [fendley2019free], and involves another set of ancillary free fermions commuting with the physical ones.
@@ V.3 The ancillary fermions
$\mathbb{E}_{2m}^{\pm}(u)$ $=\mathbb{E}_{2m-2}^{\pm}(u)+u\penalty 10000\ \mathbb{O}_{2m-2}^{\mp}(u)a_{2m}^{2}$ (50a)
$\mathbb{O}_{2m}^{\mp}(u)$ $=\mathbb{O}_{2m-2}^{\mp}(u)+u\penalty 10000\ \mathbb{E}_{2m-4}^{\pm}(u)a_{2m-1}^{2}$ (50b)
$\mathbb{O}_{2m+1}^{\pm}(u)$ $=\mathbb{O}_{2m-1}^{\pm}(u)+u\penalty 10000\ \mathbb{E}_{2m-1}^{\mp}(u)a_{2m+1}^{2}$ (50c)
$\mathbb{E}_{2m+1}^{\mp}(u)$ $=\mathbb{E}_{2m-1}^{\mp}(u)+u\penalty 10000\ \mathbb{O}_{2m-3}^{\pm}(u)a_{2m}^{2}\,.$ (50d)
@@
For a given value of $M$, let us focus on the generating function of even-order commuting charges in the algebra $\mathcal{A}_{F^{\prime}}$, namely $\mathcal{E}_{M}^{+}(u)=\mathcal{E}_{M+1}^{+}(u)$ for $M$ even (resp. $\mathcal{O}_{M}^{+}(u)=\mathcal{O}_{M+1}^{+}(u)$ for $M$ odd), and on the corresponding generating polynomial $\mathcal{P}_{M}(u)\equiv\mathbb{E}_{M}^{+}(u)$ for $M$ even (resp. $\mathbb{O}_{M}^{+}(u)$ for $M$ odd). Based on the recursion relations (50), it is easy to see that the degree of this polynomial is $2S^{\prime}=2\lfloor\frac{M+2}{6}\rfloor$, with $S^{\prime}$ pairs of opposite roots. Calling the latter $\{\pm u^{\prime}_{k}\}_{k=1,\ldots S^{\prime}}\equiv\{\pm 1/\epsilon^{\prime}_{k}\}_{k=1,\ldots S^{\prime}}$, it is then straightforward to recast the eigenvalues of the commuting charges $E^{(2n)}$ (resp. $O^{(2n)}$) in terms of the parameters $\epsilon^{\prime}_{k}$. In particular, $E^{(2)}$ (resp. $O^{(2)}$) has $2^{S^{\prime}}$ distinct eigenvalues of the form
@@
Going further, we can construct the corresponding fermionic creation and annihilation operators. Since the derivation is similar to those of Ref. [fendley2019free], we omit it and only provide the final result: the ancillary creation and annihilation operators read
$\Psi^{\prime}_{\pm k}\propto\begin{cases}\mathcal{E}_{M}(\sqrt{\pm u_{k}})\penalty 10000\ E\penalty 10000\ \mathcal{E}_{M}(\sqrt{\mp u_{k}})\qquad\mbox{for $M$ even}\\ \mathcal{O}_{M}(\sqrt{\pm u_{k}})\penalty 10000\ O\penalty 10000\ \mathcal{O}_{M}(\sqrt{\mp u_{k}})\qquad\mbox{for $M$ odd}\end{cases}\,$, (52)
@@
V.4 The symmetry algebra $\mathcal{A}_{F^{\prime}}$ in terms of the ancillary fermions
@@
We are therefore ready to state our final proposition, namely that the algebra $\mathcal{A}_{F^{\prime}}$ is generated by all bilinears in the ancillary fermions $\Psi^{\prime}_{k}$. This is confirmed by a counting of dimensions presented in Table 1, cf. Sec. III, where it is checked that the ancillary fermionic modes, together with the algebra $\mathcal{A}_{\widetilde{D}}$ of Sec. IV, indeed accounts for all the degeneracies observed in the spectrum.
V.5 The remaining zero mode for $M=6k+3$
As mentioned above, for the special case $M=6k+3$, the algebras $\mathcal{A}_{\widetilde{D}}$ and $\mathcal{A}_{F^{\prime}}$ are enough to lift all degeneracy factors of the spectrum, but one. This can be explained by the existence of an additional zero mode, which commutes both with the physical fermions and with $\mathcal{A}_{\widetilde{D}}$ and $\mathcal{A}_{F^{\prime}}$. For $M=3$, such an operator can easily be found as $(h_{1}+h_{2})\chi_{M}$. For $M=9$, it can be found as a bilinear in the $\{h_{i}\}_{i\leq M-1}$ multiplied by $\chi_{M}$. For general $M=6k+3$ we conjecture that an operator of such a form should exist, namely some multilinear combination of $\{h_{i}\}_{i\leq M-1}$ multiplied by $\chi_{M}$. Furthermore, it should be possible to relate the existence of this operator to the zero modes evidenced in [vona2014exact] more generally for all $M\notin 3\mathbb{N}+1$. However, a difference between our approach and the zero modes of [vona2014exact] is that the latter do not necessarily commute with the rest of the symmetry algebra, that is, with $\mathcal{A}_{\widetilde{D}}$ and $\mathcal{A}_{F^{\prime}}$.
VI Outlook
@@
Our work opens several natural directions. Most naturally, our results provide new tools for the computation of correlations functions, both in an out of equilibrium. Previously, this problem was tackled by seeking a solution to the so-called “inverse problem”, consisting in expressing local spin operators in terms of fermionic ones. In particular Ref. [vona2014exact] found a few special local spin operators that admit a simple representation in terms of the free fermionic operators, making it possible to study their dynamics explicitly. In our work, we have completely characterized the set of operators generating the symmetry algebra, which is supported on the degenerate subspace and, by definition, not evolving under the FFD Hamiltonian dynamics. This allows us go further, and consider spin operators admitting a simple representation not only in terms of the fermionic ones, but also of the elements of the symmetry algebra. Finally, another natural direction pertains to extending our constructions to different FFD models, including the one presented in Ref. [fendley2024free], the parafermionic generalizations developed in Refs. [alcaraz2020free, alcaraz2020integrable], the FFD quantum circuits studied in Ref. [fukai2025quantum] and more generally the families FFD-solvable models characterized by their frustration graphs [fukai2025free, chapman2023unified]. The work [chapman2023unified], in particular, introduces a further fragmentation of the Hilbert space in terms of cycle symmetries: it would be interesting to understand how this relates to the Hilbert space factorization presented in this work. These directions will be explored in future research.